### v3

@@ 6.2 Polychoric correlation matrix
The two simulation studies demonstrate that already a small degree of partial misspecificaion to uninformative responses, such as careless responses, can render the commonly employed ML estimator unreliable, while the proposed robust estimator retains good accuracy even in the presence of a considerable number of uninformative responses. On the other hand, when the model is correctly specified, the ML estimator and robust estimator produce equivalent estimates.
@@ 7.1 Background and study design
We demonstrate our proposed method on empirical data by using a subset of the 100 unipolar markers of the Big Five personality traits (Goldberg,, 1992). Each marker is a an item comprising a single English adjective (such as “bold” or “timid”) asking respondents to indicate how accurately the adjective describes their personality using a 5-point Likert-type rating scale (very inaccurate, moderately inaccurate, neither accurate nor inaccurate, moderately accurate, and very accurate). Here, each Big Five personality trait is measured with six pairs of adjectives that are polar opposites to one another (such as “talkative” vs. “silent”), that is, twelve items in total for each trait. It seems implausible that an attentive respondent would choose to agree (or disagree) to both items in a pair of polar opposite adjectives. Consequently, one would expect a strongly negative correlation between polar adjectives if all respondents respond attentively (Arias et al.,, 2020).
@@
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^5] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. Furthermore, the data of Arias et al., (2020) have been made publicly available by the authors at https://osf.io/n6krb. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
@@
Due to the suspected presence of careless respondents, we apply our proposed method to estimate the polychoric correlation coefficients between all $\binom{12}{2}=66$ unique item pairs in the neuroticism scale to obtain an estimate of the scale’s (polychoric) correlation matrix. The results of the remaining two scales are qualitatively similar and are reported in Appendix C. We estimate the polychoric correlation matrix twice: via classic maximum likelihood and via our proposed robust alternative, with tuning parameter $c=0.6$. The results remain qualitatively similar for different finite choices of $c$.
@@ 7.2 Results
Figure 6: Difference between absolute estimates for the polychoric correlation coefficient of our robust estimator and the MLE for each item pair in the neuroticism scale, using the data of Arias et al., (2020). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite. The individual estimates of each method are provided in Table C.2 in Appendix C.
Sample cor ML Robust
Parameter Estimate SE Estimate SE Estimate SE
$\rho$ $-0.562$ 0.031 $-0.618$ 0.025 $-0.925$ 0.062
$a_{1}$ $-1.370$ 0.061 $-1.570$ 0.276
$a_{2}$ $-0.476$ 0.043 $-0.560$ 0.203
$a_{3}$ $0.121$ 0.042 $0.109$ 0.187
$a_{4}$ $1.060$ 0.054 $1.080$ 0.105
$b_{1}$ $-0.857$ 0.049 $-0.905$ 0.073
$b_{2}$ $-0.004$ 0.041 $-0.040$ 0.091
$b_{3}$ $0.608$ 0.045 $0.640$ 0.364
$b_{4}$ $1.580$ 0.071 $1.171$ 0.811
Table 3: Parameter estimates with standard errors (SEs) for the correlation between the neuroticism adjective pair “envious” and “not envious” in the data of Arias et al., (2020), using Pearson sample correlation, ML, and robust estimator with tuning constant $c=0.6$. Each adjective item has five Likert-type answer categories. The sample correlation coefficient does not model thresholds, hence no estimates can be reported for them.
@@
Figure 6 visualizes the difference in absolute estimates for the polychoric correlation coefficient between all 66 unique item pairs in the neuroticism scale. For all unique pairs, our method estimates a stronger correlation coefficient than ML. The differences in absolute estimates on average amount to 0.083, ranging from only marginally larger than zero to a substantive 0.314. For correlations between polar opposite adjectives, the average absolute difference between our robust method and ML is 0.151. The fact that a robust method consistently yields stronger correlation estimates than ML, particularly between polar opposite adjectives, is indicative of the presence of negative leverage points, which drag negative correlation estimates towards zero, that is, they attenuate the estimated strength of correlation. Here, such negative leverage points could be the responses of careless respondents who report agreement or disagreement to both items in item pairs that are designed to be negatively correlated. For instance, recall that it is implausible that an attentive respondent would choose to agree (or disagree) to both adjectives in the pair “envious” and “not envious” (cf., Arias et al.,, 2020). If sufficiently many such respondents are present, then the presumably strongly negative correlation between these two opposite adjectives will be estimated to be weaker than it actually is.
@@
To further investigate the presence of careless respondents who attenuate correlational estimates, we study in detail the adjective pair “not envious” and “envious”, which featured the largest discrepancy between the ML estimate and robust estimate in Figure 6, with an absolute difference of 0.314. The results of the two estimators and, for completeness, the sample correlation, are summarized in Table 3. The ML estimate of $-0.618$ and sample correlation estimate of $-0.562$ for the (polychoric) correlation coefficient seem remarkably weak considering that the two adjectives in question are polar opposites. In contrast, the robust correlation estimate is given by $-0.925$, which seems much more in line with what one would expect if all participants responded accurately and attentively (cf., Arias et al.,, 2020).
$X$\$Y$ 1 2 3 4 5
@@
$X$\$Y$ 1 2 3 4 5
1 $>1,000$ $>1,000$ 10.82 0.14 $-0.35$
2 $>1,000$ 9.07 $-0.20$ $-0.10$ 0.42
3 4.48 $-0.35$ $-0.01$ $-0.20$ 76.14
4 $-0.12$ $-0.08$ $-0.39$ 10.66 $>1,000$
5 $-0.11$ $-0.12$ 35.01 $>1,000$ $>1,000$
(c) PRs $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})-1$
@@
Table 4: Empirical relative frequency (top), estimated response probability (center), and Pearson residual (bottom) of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of the neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=0.6$. The complete PR values are provided in Table C.5 in the appendix.
@@
To study the potential presence of careless responses in each contingency table cell $(x,y)$ for item pair “envious” and “not envious”, Table 4 lists the PRs at the robust estimate, alongside the associated model probabilities and empirical relative frequencies. [^6] A visualization of Table 4 is provided in Figure C.5 in the appendix. A total of six cells have extremely large PR values of higher than 1,000, and, in addition, five cells have a PR of higher than 9, and one cell has a PR of higher than 4. Such PR values are far away from ideal value 0 at which the model would fit perfectly, thereby indicating a poor fit of the polychoric model for such responses. It stands out that all such poorly fitted cells are those whose responses might be viewed as inconsistent. Indeed, response cells $(x,y)=(1,1),(1,2),(2,1),(2,2)$ indicate that a participant reports that neither “envious” nor “not envious” characterizes them accurately, which are mutually contradicting responses, while for response cells $(x,y)=(4,4),(4,5),(5,4),(5,5)$ both adjectives characterize them accurately, which is again contradicting. As discussed previously, such responses might due to careless responding. The robust estimator suggests that such responses cannot be fitted well by the polychoric model and subsequently downweighs their influence in the estimation procedure by mapping their Pearson residual with the linear part of the discrepancy function $\varphi(\cdot)$ in (11). Notably, also cells $(x,y)=(1,3),(3,1),(3,5),(5,3)$ are poorly fitted. These responses report (dis)agreement to one opposite adjective, while being neutral about the other opposite. It is beyond the scope of this paper to assess whether such response patterns are also indicative of careless responding, but the robust estimator suggests that such responses at least cannot be fitted well by the polychoric model with the data of Arias et al., (2020).
@@
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of ML, amounting to about $-0.62$, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about $-0.93$ and also help identify them through unreasonably large PR values.
@@ 8 Estimation under distributional misspecification
8.1 Preliminaries
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient. Estimators for such situations where the population distribution $G$ is nonnormal have been proposed by Lyhagen & Ornstein, (2023) and Jin & Yang-Wallentin, (2017).
Huber & Ronchetti, (2009, p. 4) note that, although conceptually distinct, robustness to distributional misspecification and partial misspecification are “practically synonymous notions”. Hence, despite distributional misspecification not being covered by the misspecified distribution (7) under which we study the theoretical properties of our proposed estimator, our estimator could still offer a gain in robustness compared to ML estimation under distributional misspecification. Specifically, the robust estimator may still be useful if the central part of the nonnormal distribution $G$ is not too different from a standard bivariate normal distribution: Intuitively, if the difference between $G$ and a standard normal distribution is mainly in the tails, $G$ can be approximated by a mixture of a standard normal distribution (to cover the central part) and some other distribution $H$ (to cover the tails). The ML estimator tries to treat influential observations from the tails—which cannot be fitted well by the polychoric model—as if they were normally distributed, resulting in a possibly large estimation bias. In contrast, the robust estimator uses the normal distribution only for observations from the central part—which may fit the polychoric model well enough—and downweights observations from the tails. Thus, as long as such a mixture is a decent approximation of the nonnormal distribution $G$, the robust estimator should perform reasonably well. However, if $G$ cannot be approximated by such a mixture, neither the ML estimator nor our estimator can be expected to perform well. Overall, though, our estimator could offer an improvement in terms of robustness to distributional misspecification.

### v4

@@ 6.3 Discussion and additional simulation experiments
The two simulation studies above demonstrate that already a small degree of partial misspecification due to uninformative responses, such as careless responses, can render the commonly employed MLE unreliable, while the proposed robust estimator retains good accuracy and coverage even in the presence of a considerable number of uninformative responses. On the other hand, when the polychoric model is correctly specified, the MLE and the robust estimator produce equivalent estimates.
To further evaluate our robust estimator and investigate its limitations, we conduct additional simulation experiments in Appendix E.
The first experiment, described in Appendix E.2, is a generalization of the design in Section 6.1 with different mean shifts in the contamination distribution $H$. For small mean shifts, the proposed estimator does not improve upon the MLE, but the bias of both estimators remains reasonable. The larger the mean shift, the larger the detrimental effect on the MLE and the higher the robustness gain of our proposed estimator.
The second experiment, described in Appendix E.3, focuses on correlation shifts in the contamination distribution $H$. Specifically, the contamination distribution is the same as the true model distribution except for a sign-flipped correlation coefficient $-\rho_{*}$. For moderate correlation $\rho_{*}$, the proposed estimator does not improve upon the MLE due to substantial overlap between the true model distribution and the contamination. However, the gain in robustness increases substantially for higher correlation coefficients $\rho_{*}$.
@@ 7 Empirical application
We now demonstrate our proposed method on empirical data by using a subset of the 100 unipolar markers of the Big Five personality traits (Goldberg,, 1992).
@@ 7.1 Background and study design
Each marker is a an item comprising a single English adjective (such as “bold” or “timid”) asking respondents to indicate how accurately the adjective describes their personality using a 5-point Likert-type rating scale (very inaccurate, moderately inaccurate, neither accurate nor inaccurate, moderately accurate, and very accurate). Here, each Big Five personality trait is measured with six pairs of adjectives that are polar opposites to one another (such as “talkative” vs. “silent”), that is, twelve items in total for each trait. It seems implausible that an attentive respondent would choose to agree (or disagree) to both items in a pair of polar opposite adjectives. Consequently, one would expect a strongly negative correlation between polar adjectives if all respondents respond attentively (Arias et al.,, 2020).
@@
Arias et al., (2020) collect measurements of three Big Five traits in this way, namely extroversion, neuroticism, and conscientiousness. [^13] Arias et al., (2020) synonymously refer to neuroticism as emotional stability. Furthermore, in addition to the three listed traits, Arias et al., (2020) collect measurements of the trait dispositional optimism by using a different instrument, and another scale that is designed to not measure any construct. We do not consider these scales in this empirical demonstration. Furthermore, Arias et al., (2020) have made their data publicly available at https://osf.io/n6krb. The sample that we shall use, Sample 1 in Arias et al., (2020), consists of $N=725$ online respondents who are all U.S. citizens, native English speakers, and tend to have relatively high levels of reported education (about 90% report to hold an undergraduate or higher degree). Concerned about respondent inattention in their data, Arias et al., (2020) construct a factor mixture model for detecting inattentive/careless participants. Their model crucially relies on response inconsistencies to polar opposite adjectives and is designed to primarily detect careless straightlining responding. They find that careless responding is a sizable problem in their data. Their model finds evidence of straightliners, and the authors conclude that if unaccounted for, they can substantially deteriorate the fit of theoretical models, produce spurious variance, and overall jeopardize the validity of research results.
@@
Due to the suspected presence of careless respondents, we apply our proposed method to estimate the polychoric correlation coefficients between all $\binom{12}{2}=66$ unique item pairs in the neuroticism scale to obtain an estimate of the scale’s (polychoric) correlation matrix. The results of the remaining two scales are qualitatively similar and are reported in Appendix F. We estimate the polychoric correlation matrix via the MLE and via our proposed robust alternative with tuning parameter $c=0.6$. As a robustness check, we further investigate the effect of the choice of $c$.
@@ 7.2 Results
Figure 6: Difference between absolute estimates for the polychoric correlation coefficient of the robust estimator with $c=0.6$ and the MLE for each item pair in the neuroticism scale, using the data of Arias et al., (2020). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite. The individual estimates of each method are provided in Table F.2 in Appendix F.
@@
Figure 6 visualizes the difference in absolute estimates for the polychoric correlation coefficient between all 66 unique item pairs in the neuroticism scale. For all unique pairs, our method estimates a stronger correlation coefficient than the MLE. The differences in absolute estimates on average amount to 0.083, ranging from only marginally larger than zero to a substantive 0.314. For correlations between polar opposite adjectives, the average absolute difference between our robust method and the MLE is 0.151. The fact that a robust method consistently yields stronger correlation estimates than the MLE, particularly between polar opposite adjectives, is indicative of the presence of leverage points, which drag negative correlation estimates towards zero, that is, they attenuate the estimated strength of correlation. Here, such leverage points could be the responses of careless respondents who report agreement or disagreement to both items in item pairs that are designed to be negatively correlated. For instance, recall that it is implausible that an attentive respondent would choose to agree (or disagree) to both adjectives in the pair “envious” and “not envious” (cf., Arias et al.,, 2020). If sufficiently many such respondents are present, then the presumably strongly negative correlation between these two opposite adjectives will be estimated to be weaker than it actually is.
Robust ML Sample cor
Parameter Estimate $\widehat{\text{SE}}$ Estimate $\widehat{\text{SE}}$ Estimate $\widehat{\text{SE}}$
$\rho$ 0.925-0.925- 0.925 0.062 0.618-0.618- 0.618 0.025 0.562-0.562- 0.562 0.031
$a_{1}$ 1.567-1.567- 1.567 0.276 1.373-1.373- 1.373 0.061
$a_{2}$ 0.560-0.560- 0.560 0.203 0.476-0.476- 0.476 0.043
$a_{3}$ 0.1100.1100.110 0.187 0.1210.1210.121 0.042
$a_{4}$ 1.0761.0761.076 0.105 1.0591.0591.059 0.054
$b_{1}$ 0.905-0.905- 0.905 0.073 0.857-0.857- 0.857 0.049
$b_{2}$ 0.040-0.040- 0.040 0.091 0.004-0.004- 0.004 0.041
$b_{3}$ 0.6400.6400.640 0.364 0.6080.6080.608 0.044
$b_{4}$ 1.1711.1711.171 0.811 1.5831.5831.583 0.071
Table 3: Parameter estimates and standard error estimates ($\widehat{\text{SE}}$) for the correlation between the neuroticism adjective pair “not envious” and “envious” in the data of Arias et al., (2020), using the robust estimator with $c=0.6$, the MLE, and the Pearson sample correlation. Each adjective item has five answer categories. Note that the Pearson sample correlation does not model thresholds.
@@
To further investigate the presence of careless respondents who attenuate correlational estimates, we study in detail the adjective pair “not envious” and “envious”, which featured the largest discrepancy between the ML estimate and the robust estimate in Figure 6, with an absolute difference of 0.314. The results of the two estimators and, for completeness, the sample correlation, are summarized in Table 3. The ML estimate of 0.618-0.618- 0.618 and sample correlation estimate of 0.562-0.562- 0.562 for the (polychoric) correlation coefficient seem remarkably weak considering that the two adjectives in question are polar opposites. In contrast, the robust correlation estimate is given by 0.925-0.925- 0.925, which seems much more in line with what one would expect if all participants responded accurately and attentively (cf., Arias et al.,, 2020).
[…]
@@
[…]
1 >1,000>1,000> 1 , 000 >1,000>1,000> 1 , 000 10.81 0.14 0.35-0.35- 0.35
2 >1,000>1,000> 1 , 000 9.06 0.20-0.20- 0.20 0.10-0.10- 0.10 0.42
3 4.48 0.35-0.35- 0.35 0.01-0.01- 0.01 0.20-0.20- 0.20 76.11
4 0.12-0.12- 0.12 0.08-0.08- 0.08 0.39-0.39- 0.39 11.66 >1,000>1,000> 1 , 000
5 0.11-0.11- 0.11 0.12-0.12- 0.12 34.98 >1,000>1,000> 1 , 000 >1,000>1,000> 1 , 000
(c) Pearson residuals $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})-1$
@@
Table 4: Empirical relative frequency (top), estimated response probability (center), and Pearson residual (PR) (bottom) of each response $(x,y)$ for the item pair “not envious” ($X$) and “envious” ($Y$) in the measurements of Arias et al., (2020) of the neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed via the robust estimator with tuning constant $c=0.6$. The complete PR values are provided in Table F.5 in the appendix.
@@
To study the potential presence of careless responses in each contingency table cell $(x,y)$ for item pair “not envious” and “envious”, Table 4 lists the PRs at the robust estimate, alongside the associated model probabilities and empirical relative frequencies. [^14] A visualization of Table 4 is provided in Figure F.4 in the appendix. A total of six cells have extremely large PR values of higher than 1,000, and, in addition, five cells have a PR of higher than 9, and one cell has a PR of higher than 4. Such PR values are far away from ideal value 0 at which the model would fit perfectly, thereby indicating a poor fit of the polychoric model for such responses. It stands out that all such poorly fitted cells are those whose responses might be viewed as inconsistent. Indeed, response cells $(x,y)\in\{(1,1),(1,2),(2,1),(2,2)\}$ indicate that a participant reports that neither “not envious” nor “envious” characterizes them accurately, which are mutually contradicting responses, while for response cells $(x,y)\in\{(4,4),(4,5),(5,4),(5,5)\}$ both adjectives characterize them accurately, which is again contradicting. As discussed previously, such responses might be due to careless responding. The robust estimator suggests that such responses cannot be fitted well by the polychoric model and subsequently downweighs their influence in the estimation procedure by mapping their Pearson residual with the linear part of the discrepancy function $\varphi(\cdot)$ in (11). Notably, also cells $(x,y)\in\{(1,3),(3,1),(3,5),(5,3)\}$ are poorly fitted. These responses report (dis)agreement to one opposite adjective, while being neutral about the other opposite. It is beyond the scope of this paper to assess whether such response patterns are also indicative of careless responding, but the robust estimator suggests that such responses at least cannot be fitted well by the polychoric model with the data of Arias et al., (2020).
As a robustness check on the role of tuning constant $c$, Figure 7 visualizes the point estimate $\widehat{\rho}_{N}$ of the polychoric correlation between the item pair “not envious” and “envious” for various values of $c$. The point estimate stays relatively constant for $c$ between 0 and 0.75, with $\widehat{\rho}_{N}$ ranging between 0.95-0.95- 0.95 and 0.92-0.92- 0.92. Just after $c=0.75$, $\widehat{\rho}_{N}$ abruptly jumps to about 0.85-0.85- 0.85, before it stabilizes again and slowly transitions to the ML estimate of 0.62-0.62- 0.62 (see Table 3) for $c\rightarrow\infty$. Since $\widehat{\rho}_{N}$ very slowly approaches the value of the ML estimate, we only visualize choices of $c$ up to 2 in Figure 7. The instability of the estimate around $c=0.75$ suggests that $c$ should be chosen below this value to avoid disproportionate influence of poorly fitting cells. For the broad range of $c\leq 0.75$, we obtain a robust finding of “not envious” and “envious” having a very strong negative correlation after accounting for likely careless responding.
Figure 7: Estimates of the polychoric correlation between the items “not envious” and “envious” in the data of Arias et al., (2020) for various choices of the tuning constant $c$ ($x$-axis). The dashed vertical line marks the default value of $c=0.6$.
@@
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of the MLE, amounting to about 0.62-0.62- 0.62, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about 0.93-0.93- 0.93 and also help identify them through unreasonably large PR values. On a final note, our findings from the empirical application align remarkably well with those of the simulation from Appendix E.3, therefore strenghening their validity. We provide a detailed discussion of this similarity in Appendix E.3.3.
@@ 8 Estimation under distributional misspecification
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient. Estimators for such situations where the population distribution $G$ is nonnormal have been proposed by Lyhagen & Ornstein, (2023), Jin & Yang-Wallentin, (2017), and Roscino & Pollice, (2006).
8.1 Distributional vs. partial misspecification
Huber & Ronchetti, (2009, p. 4) note that, although conceptually distinct, robustness to distributional misspecification and partial misspecification are “practically synonymous notions”. Hence, despite distributional misspecification not being covered by the partial misspecification framework under which we study the theoretical properties of our proposed estimator, our estimator could still offer a gain in robustness compared to the MLE of the polychoric model under distributional misspecification. Specifically, the robust estimator may still be useful if the central part of the nonnormal distribution $G$ is not too different from a standard bivariate normal distribution. Intuitively, if the difference between $G$ and a standard normal distribution is mainly in the tails, $G$ can be approximated by a contaminated distribution as in (7), with the standard normal distribution covering the central part and some contamination distribution $H$ covering the tails. The polychoric MLE tries to treat influential observations from the tails—which cannot be fitted well by the polychoric model—as if they were normally distributed, resulting in a possibly large estimation bias. In contrast, the robust estimator uses the normal distribution only for observations from the central part—which may fit the polychoric model well enough—and downweights observations from the tails. Thus, as long as such a contaminated normal distribution is a decent approximation of the nonnormal distribution $G$, the robust estimator should perform reasonably well. However, if $G$ cannot be approximated by such a contaminated normal distribution, neither the polychoric MLE nor our estimator can be expected to perform well. Overall, though, our estimator could offer an improvement in terms of robustness to distributional misspecification.